# SRLM: Human-in-Loop Interactive Social Robot Navigation with Large Language Model and Deep Reinforcement Learning

### v2

@@ Abstract
An interactive social robotic assistant must provide services in complex and crowded spaces while adapting its behavior based on real-time human language commands or feedback. In this paper, we propose a novel hybrid approach called Social Robot Planner (SRLM), which integrates Large Language Models (LLM) and Deep Reinforcement Learning (DRL) to navigate through human-filled public spaces and provide multiple social services. SRLM infers global planning from human-in-loop commands in real-time, and encodes social information into a LLM-based large navigation model (LNM) for low-level motion execution. Moreover, a DRL-based planner is designed to maintain benchmarking performance, which is blended with LNM by a large feedback model (LFM) to address the instability of current text and LLM-driven LNM. Finally, SRLM demonstrates outstanding performance in extensive experiments. More details about this work are available at: https://sites.google.com/view/navi-srlm.
@@ I Introduction
Navigating in the human-filled spaces is a crucial aspect of social robots to support various of advanced services, such as cooperating or walking together with users. The socially-aware navigation (SAN) task faces two main challenges: the aspect of real-time highly volatile user requests or feelings, and the constraint of managing socially compliant or acceptable navigation behaviors within dynamic environments. With the developments in robotics and artificial intelligence technologies, current approaches have addressed the aforementioned issues to implement social robots in public environments [1]. These approaches are inspired by significant insights from fields such as machine learning [2], sociology [3], analytical mechanics [4], algebra and geometry [5], and others.
However, both existing learning-based and conventional approaches exhibit limited adaptability when it comes to real-time response requirements. Therefore, we design an interactive Social Robot navigation Large Model (SRLM) that can infer and execute users’ real-time commands, leveraging the promising potential of large language model (LLM) in human language understanding. For instance, users can adjust robot configuration or behavioral styles corresponding to personal feelings in real-time, a challenge that current state-of-the-art (SOTA) planners struggle with due to the fixed parameters of converged policies. Hence, SRLM interprets real-time human feedback inference via an LLM as high-level global information to guide low-level actions execution. This interactive framework not only enhances user experience but also boosts performance, as shown in Fig. 1.
Figure 1: An illustration of human-in-loop interactive social robot navigation
Despite the significant applications of LLM in a wide range of areas, such as robotics [6], HRI [7], logical reasoning [8, 9], etc, the deployment of LLM-based social navigation planners is still limited by the gaps in navigation information (such as location and velocity) to language text and LLM’s training dataset. Moreover, most recent LLM-based navigation systems [6, 10] directly generate macro-action functions rather than real low-level motion control instructions, due to the insensitivity and misunderstanding of LLM regarding continuous space’s numerical values. To adapt the ability of large models in navigation aspects, SRLM designs a textual embedding encoder and a social navigation prompt to convert environmental information for generating low-level navigation actions directly.
Our goal is to develop an interactive social navigation artificial general intelligence aligned with each user’s personal preferences. Additionally, we organically incorporate existing deep reinforcement learning-based SOTA navigation planners with LLM-based approaches into a large navigation model. This consideration arises from the challenge that recent LLM-based planners face in handling complex dynamic optimization problems in crowded environments. The main contributions of this paper can be summarized as follows:
@@
  • We propose SRLM, a human-in-loop interactive social robot navigation framework driven by LLMs and deep reinforcement learning (DRL). SRLM can execute personalized social robot tasks according to users’ requests, preferences, and real-time feedback from human language, serving as an interactive social robot assistant.
  • SRLM leverages an advanced DRL-based social planner and a language navigation model to generate socially compliant robot behaviors. Additionally, the language navigation model (LNM) memory mechanism can store temporal data and provide long-term evaluations and feedback to refine DRL-based and LLM-based social planners.
  • SRLM adapts the properties of user tasks and personal preferences driven by LLM as high-level global guidance into reinforcement learning from human feedback (RLHF) modifications. In this setup, the reward network from RLHF can be further aligned by users rather than human training supervisors in the DRL-based planner.
@@
  • SRLM demonstrates robust and promising exhibitions of socially compliant behaviors in various experiments.
@@ II Background
LLM-driven Navigation. Inspired by the promising of LLM across a wide range of applications, successes driven by extensive computational resources and advanced machine learning methods have motivated massive research in robotics and HRI. For instance, [11] designs an LLM-driven mobile manipulator that offers services or infers user preferences from human language requests. The inference ability of LLM is essential for interactive social robots capable of tracking real-time user language feedback to adjust robot behaviors. Here, we leverage a high-level LLM block as an interactive framework to generate global guidance in response to human feedback.
Moreover, the context semantic inference of LLM can also be employed for navigation execution. For instance, [12] introduces LLM-guide visual language navigation, where planners are controlled by conclusions drawn from textual robot perception features provided by LLM. However, despite the reasonable evidences and dependencies for LLM in similar social navigation environments, the data schema and inference difficulty of social navigation are more complicated than those of visual or language navigation. Therefore, we abstract social navigation environmental features into textual data to adapt an LLM-driven large navigation model capable of inferring social interaction and directly generating low-level robot actions.
To further improve the inference ability, the chain-of-thoughts [8] technology has been developed to generate intermediate steps of inference process. The chain-type construction provides more generative information and an adjustable method. Additionally, structures such as trees and graphs have been proposed for a better inference structure, as seen in [9, 13]. Thus, considering the ephemeral limitations of LLM inference, particularly due to potential probabilistic illusions in LLM reasoning and the insensitivity of sequential floating-point numbers in LLM, the reinforcement learning navigation model (RLNM) is introduced to maintain baseline performance. This performance is adaptively fused by the large feedback model (LFM) via a Graph-of Thoughts (GoT) construction.
Socially Aware Robot Navigation. After early applications of robotics in social navigation society, such as MINERVA [14], socially aware robot navigation tasks are primarily conducted via decoupled and coupled strategies [15]. Decoupled approaches infer pedestrian motion intents and patterns to construct potential safe areas for planing. However, the separation of modeling and planning often overlooks potential cooperation, leading to the establishment of limited feasible spaces, known as the freezing robot problem [16], particularly with the increasing presence of humans. Alternatively, coupled approaches encode potential cooperation into navigation inference to address unwarranted ignorance.
@@
On the other hand, explicitly coupled approaches are implemented through game theoretic planning [17], Gaussian processes [18], and topology analysis [3]. However, the challenge of optimization in highly dynamic environments restricts the further deployment of conventional approaches [19] and explicitly coupled approaches, especially with the increasing complexity of the environment. Recently, the paradigm of cooperative collision avoidance has facilitated a set of promising works [2, 20], which implicitly approximate human-like navigation awareness and insight through advanced neural networks to encode potential human-robot cooperation and compliance with social norms into robot behavior. These neural networks are then trained using DRLs to iterate through different situations. For instance, efforts have been made in the development of neural network technologies for social navigation, such as attention mechanisms [21], graph construction [20], and transformers [2].
@@
Despite aforementioned neural networks being utilized to evaluate underlying human-robot interaction and pedestrian intents for cooperative collision avoidance, human preferences are still not well represented. [22] incorporates the high-order uncertainty of human movements as pedestrian preference distributions into the social navigation planner based on variational analysis. Furthermore, the SAN task is also motivated by the direct involvement of human intelligence through RLHF in [2, 23], where human expectations and social norms are studied and embedded by a reward neural network to train the policy. More currently, [24] extends the learning-based planner in multi-robot scenarios.
Learning-based approaches generally exhibit benchmarking efficacy. However, the converged DRL policies often result in degradation, with limited generalization ability in unfamiliar scenes, and they are also difficult to adapt to real-time user feelings. Hence, we leverage LNM and LFM to improve the adaptability and robustness of DRL-base planners, incorporating the inference capability of large models for both real-time human feedback and low-level motion execution.
@@ III Preliminary
SRLM is an interactive social navigation framework that understands user commands (e.g. “Pick up my bag to me") and personal preferences (such as the user needing a larger privacy area with the robot) into a set of high-level instructions as global guidance from human language input. Herein, SRLM blends a high-level and low-level execution system, in which task objectives, such as point-to-point (P2P), human-guide (HG), human-follow (HF), user preferences (privacy distance), and social norm property (whether to wait each pedestrian), are composed into high-level global guidance for low-level robot action generation.
@@
Subsequently, the pre-trained LNM and RLNM generate low-level robot actions with respect to textual or featurized HRI state presentations and the above global guidances. Moreover, the global guidance information can be modified or updated by real-time human feedback as well, such as when a user adds personal feelings or changes task objectives, prompting instructions to be replanned. Lastly, an additional feedback and memory mechanism is adapted to adjust LNM & RLNM incorporating behaviors from past trajectories. The real-time evaluation and feedback mechanism provide an adaptive heuristic to take their advantages. For instance, the DRL-based action execution model maintains a lower bound when the LM-based model encounters explicit mistakes or dangers, thus providing fundamental performance as the social navigation benchmark. On the other hand, the LM-based action execution model incorporates user personal modifications from user language to improve the robustness of DRL-based action execution model and RLHF reward network.
Figure 2: SRLM architecture: SRLM is implemented as a human-in-loop interactive social robot navigation framework, which executes human commands based on LM-based planner, feedback-based planner, and DRL-based planner incorporating. Firstly, users’ requests or real-time feedbacks are processed or replanned to high-level task guidance for three action executors via LLM. Then, the image-to-text encoder and spatio-temporal graph HRI encoder convert robot local observation information to features as LNM and RLNM input, which generate RL-based action, LM-based action, and feedback-based action. Lastly, the above three actions are adaptively fused by a low-level execution decoder as the robot behavior output of SRLM.
@@
Herein, the interactive social robot navigation problem is formulated as a Dec-POSMDP (decentralized particularly observable semi-Markov decision process) problem based on [24], characterized by the tuple $\langle\mathcal{S},\mathcal{U},\mathcal{A},\Omega,\mathcal{O},\mathcal{P},\mathcal{R},\mathbf{R},\mathcal{C},\mathcal{S}_{0},\gamma,\rm N,\Upsilon\rangle$. Here, $\rm\mathbf{{s}}_{\rm{t}}=[\rm\mathbf{s}^{r}_{\rm{t}},\rm\mathbf{s}^{h_{1}}_{\rm{t}},\cdots,\rm\mathbf{s}^{h_{N}}_{\rm{t}}]\in\mathcal{S}$ denotes the fully observable and unobserved states of the robot and humans at the $\rm t$-th timestep, belong to the state space, with the observable state denoted by ${\rm\mathbf{s}^{o}_{\rm{t}}}=[p_{\rm x},p_{\rm y},v_{\rm{x}},v_{\rm{y}},\rho]$ covering individual position, velocity, and radius information that can be estimated by the robot. Accordingly, pedestrians’ personal preferences and intent goals remain unobserved by robots, represented as $\mathbf{s}^{\rm{uo}}_{\rm{t}}=[{g}_{\rm{x}},{g}_{\rm{y}},v_{\rm{pref}}]$. Moreover, $\mathrm{{u}_{t}}\in\mathcal{U}$ represents robot macro-action (MA), such as waypoint locations or robot operations, which are adaptive to real-time user requirements or feedback, while robot local-action (LA) are denoted by $\mathrm{a_{t}}=[a_{x},a_{y}]\subseteq\mathcal{A}$, representing acceleration. $\mathcal{O}\rm{(s^{o}|(s,a))}$ denotes the observation probability of the robot in the observation space $\Omega$, and $\mathcal{P}$ represents the state transition probability. $\mathcal{R},\mathbf{R}$ represent the MA reward and LA reward space separately, wherein the LA reward neural network $r_{\varphi}$ is trained by the RLHF procedure. In particular, SRLM also introduces user commands as an additional large model reward $r_{LM}$ for LA reward (e.g., LLM designs a reward function term “$r_{LM}=-r_{a}|(dis_{r,i}<2m)$" corresponding to user language “Please maintain at least 2 $m$ distance to me"). The MA reward is generated by the following objective: ${\mathcal{R}}(\mathrm{{s},{u}})=\mathop{\arg\max}\mathbb{E}[\textstyle\sum_{\rm t=0}^{\rm T}\gamma^{\rm t}{\mathbf{R}}({\rm s}_{\rm t},{\rm a}_{\rm t})|{\rm a}_{\rm t}\sim{\rm u}]$. Additionally, $\mathcal{C}$ represents the conditional function, which can be updated by user language comments from LLM, $\mathcal{S}_{0}$ is the initial distribution, $\rm N$ is the number of pedestrians, $\gamma\in[0,1]$ is a discount factor, and $\Upsilon$ denotes the LLM global guidances wherein task objectives, user preferences, and robot properties are involved.
@@
The interactive social navigation task statement can be viewed as a condition of a single robot from the multi social robot navigation task definition [24], with the same robot kinematic and dynamic configurations. For further definitions and theorems, refer to [25, 24].

### v3

# Unifying Large Language Model and Deep Reinforcement Learning for Human-in-Loop Interactive Socially-aware Navigation

@@ Abstract
Navigating human-filled spaces is crucial for the interactive social robots to support advanced services, such as cooperative carrying, which enables service provision in complex and crowded environments while adapting behavior based on real-time human language commands or feedback. However, existing social robot navigation planners face two major challenges: managing real-time user inputs and ensuring socially compliant behaviors in unfamiliar, zero-shot environments. In response, we introduce SALM, an interactive, human-in-loop Socially-Aware navigation Large Language Model framework that dynamically integrates deep reinforcement learning (DRL) with large language model (LLM) capabilities. SALM leverages contextual semantic understanding from real-time human-robot interactions to convert high-level user commands into precise, low-level control actions. A high-level LLM module parses user input, guiding the simultaneous generation of navigation commands by both a large language navigation model (LNM) and a DRL-based navigation model (RLNM). A memory mechanism archives temporal data for continuous refinement, while a multi-step graph-of-thoughts inference-based large language feedback model adaptively fuses the strengths of both planning approaches. Experimental evaluations demonstrate that SALM not only enhances navigational precision in crowded, dynamic environments but also significantly improves system adaptability, offering tailored behaviors that align with individual user preferences and real-time feedback. More details and videos about this work are available at: https://sites.google.com/view/navi-salm.
@@ I Introduction
Navigating in human-filled spaces is crucial for social robots to support various of advanced services [1, 2, 3], such as human-robot cooperative carrying. The socially-aware navigation (SAN) task still faces two primary challenges: managing highly volatile, real-time user requests or sentiments, and ensuring socially compliant navigation behaviors in zero-shot unfamiliar environments. Recent developments in robotics and artificial intelligence have led to approaches that address these challenges for deploying social robots in public spaces [4, 5]. These solutions draw upon insights from machine learning [6], sociology [7], analytical mechanics [8, 9], algebra and geometry [10], among other fields, to achieve robust and socially-aware navigation.
However, both existing learning-based and conventional approaches demonstrate limited adaptability and transferability when addressing real-time response requirements in unfamiliar or zero-shot environments. In response, we introduce an interactive Socially Aware navigation large Language Model (SALM) that dynamically adapts to environmental changes by leveraging the potential of large language models (LLMs) for understanding contextual semantics in human-robot interaction (HRI) based on users’ real-time commands and feedback. For example, users can adjust the robot’s configuration or behavioral style according to their personal preferences, which a flexibility that current state-of-the-art (SOTA) planners often lack due to their fixed, converged policy parameters, as shown in Fig. 1.
Figure 1: An illustration of human-in-the-loop interactive social navigation: integrating DRL and LLM enables the social robot to adaptively respond to environmental changes such as task target and preferred social distance from real-time human feedback.
On the other hand, despite the widespread applications of LLMs in robotics [11], HRI [12], and logical reasoning [13, 14], the deployment of LLM-based social navigation planners remains constrained. These limitations arise from gaps in real-time navigation information and biases inherent in standard pre-training datasets. Moreover, many current LLM-based navigation systems [11, 15] generate macro-action functions rather than low-level motion control commands, largely due to LLMs’ insensitivity to continuous numerical values. To address these challenges and improve navigational precision, SALM integrates a textual embedding encoder with a tailored social navigation prompt. This combination effectively translates environmental data into direct low-level navigation commands, ensuring consistent inference between environmental semantic understanding and path planning.
We propose an interactive, embodied intelligent social navigation system that is tailored to each user’s personal preferences and real-time feedback. Our approach seamlessly integrates a SOTA DRL navigation planner with LLM-based techniques into a unified navigation framework. This hybrid methodology addresses the limitations of current LLM-based planners, particularly their challenges in handling complex dynamic optimization problems in crowded environments, while also enhancing the transferability of RL-based planners. The main contributions of this paper are as follows:
@@
  • We propose SALM, a human-in-loop interactive Socially-Aware navigation large Language Model framework, incorporating both LLMs and deep reinforcement learning (DRL). SALM harnesses users’ requests, preferences, and real-time feedback from human language to craft tailored tasks for social robots, functioning as an engaging, interactive assistant.
  • SALM executes interactive tasks using a high-level user parsing LLM module that directs the generation of low-level control behaviors from both the large language navigation model (LNM) and the reinforcement learning navigation model (RLNM). This approach facilitates furthermore contextual semantic understanding to produce socially compliant robot behaviors. Additionally, the memory mechanism of LNM archives temporal data, providing long-term evaluations and feedback to continuously refine the DRL-based and LLM-based social planners.
  • SALM adaptively integrates the unique strengths of both the LNM and the RLNM by incorporating a large language feedback model (LFM) scorer, utilizing multi-step Graph-of-Thoughts (GoT) inference technology.
@@
  • SALM demonstrates robust and promising exhibitions of socially compliant behaviors in various experiments.
@@ II Background
LLM-driven Navigation: Inspired by the promising potential of LLM across diverse applications, recent breakthroughs powered by extensive computational resources and advanced machine learning techniques have catalyzed significant research in robotics [8, 15] and HRI [16, 12]. The robust inference capabilities of LLM are essential for interactive social robots, substantially enhancing both the transferability and generalizability of pre-trained configurations to zero-shot environments. In this work, we leverage a high-level LLM module to parse user requirements as global guidance for low-level execution LLM as LNM.
Recent LLM-based social navigation planners [17, 18, 19, 20] have demonstrated impressive capabilities in inferring contextual semantics and planning paths in human-populated environments. However, these approaches typically generate only semantic goals (e.g., go to position $p_{A}$) rather than direct low-level control signals (e.g., [$v_{x},v_{y}$]) required for precise robot motion. This lack of low-level execution impedes both the accuracy and continuity of navigation. To address this limitation, SALM abstracts environmental features of social navigation into textual information for effectively capturing dynamic HRI nuances via the LNM block, enabling the direct generation of low-level robot velocity actions, as shown in Fig. 3.
Moreover, to address inherent limitations in LLM inference, such as potential probabilistic hallucinations in reasoning and the insensitivity to sequential floating-point numbers, the RLNM is introduced to ensure robust baseline performance. The LNM and RLNM are then adaptively integrated via the LFM, which evaluates and scores action pairs from both models to select the final action. To further enhance inference capabilities, SALM employs chain-of-thoughts [13, 14, 21] technology, which generates intermediate reasoning steps, offering richer generative insights and increased flexibility. Consequently, a graph inference structure is implemented within the LFM for logical and detailed action evaluation, as illustrated in Fig.4.
Socially Aware Robot Navigation: Building on early applications of robotics in social navigation, such as MINERVA [22], current research in socially aware robot navigation typically follows either decoupled or coupled strategies [23]. In decoupled approaches, pedestrian motion intents and patterns are first inferred to delineate potential safe areas for planning. However, this separation between modeling and planning can neglect cooperative behaviors, often resulting in overly restrictive feasible spaces and triggering the "freezing robot problem" [24] that becomes more pronounced with higher human density. In contrast, coupled approaches integrate cooperative interactions directly into the navigation inference process, thereby mitigating these limitations.
@@
Alternatively, explicitly coupled approaches are implemented through game-theoretic planning [25], Gaussian processes [26], and topology analysis [7]. However, the optimization challenges posed by highly dynamic environments hinder the practical deployment of conventional methods like ORCA [27] and other explicitly coupled techniques, especially as environmental complexity increases. Recently, the cooperative collision avoidance paradigm has spurred promising research that implicitly captures human-like navigation awareness. These approaches leverage advanced neural networks to encode potential human-robot cooperation and compliance with social norms into robot behavior. Trained using deep reinforcement learning (DRL), these networks iterate through diverse scenarios, with recent advancements incorporating technologies such as attention mechanisms [28], graph attention networks [29], transformers [6], and diffusion models [8].
@@
Despite the utilization of DRL to evaluate underlying human–robot interactions and pedestrian intents for cooperative collision avoidance, human preferences remain inadequately represented. Thus, [30] incorporates the high-order uncertainty of human movements by modeling pedestrian preference distributions within a social navigation planner based on variational analysis. Furthermore, the SAN task has been advanced by directly involving human intelligence through reinforcement learning from human feedback, as demonstrated in [6, 8], where human expectations and social norms are integrated into the training policy via a reward neural network. More recently, extensions to the learning-based SAN planner have been proposed for multi-robot scenarios [31, 32], further broadening the scope of cooperative navigation strategies.
While learning-based approaches typically perform well on benchmarks, DRL policies often struggle in zero-shot unfamiliar environments, leading to performance degradation and a disconnect from real-time human sentiment. For example, a DRL-based policy trained with a fixed social distance parameter (d = 0.4) may fail to adapt when a user interacts with the robot in ways that require different proximities such as approaching closely to hand over a package (d = 0) or keeping a greater distance (d = 1.5). To address this challenge, we integrate the large language navigation model (LNM) and the large language feedback model (LFM) into SALM. This integration harnesses the inference capabilities of large language models to process real-time human feedback and generate precise low-level motion commands, thereby enhancing the adaptability and robustness of DRL-based planners.
Figure 2: SALM architecture: SALM is implemented as a human-in-loop interactive social robot navigation framework, which executes human commands based on LM-based planner, feedback-based planner, and DRL-based planner incorporating. Firstly, users’ requests or real-time feedback are processed or replanned to high-level task guidance for three action executors via LLM. Then, the image-to-text encoder and spatio-temporal graph HRI encoder convert robot local observation information to features as LNM and RLNM input, which generate RL-based action, LM-based action, and feedback-based action. Lastly, the above three actions are adaptively fused by a low-level execution decoder as the robot behavior output of SALM.
@@ III Preliminary
SALM is an interactive social navigation framework that understands user commands (e.g., “Pick up my bag to me") and personal preferences (such as the user needing a larger privacy area with the robot) into a set of high-level instructions as global guidance from human language input. Herein, SALM blends a high-level and low-level execution system in which task objectives, such as point-to-point (P2P), human-follow (HF), user preferences (privacy distance), and social norm property (whether to wait each pedestrian), are composed into high-level global guidance for low-level robot action generation.
@@
Subsequently, the pre-trained LNM and RLNM generate low-level robot actions with respect to textual or featurized HRI state presentations and the above global guidance. Moreover, the global guidance information can be modified or updated by real-time human feedback as well, such as when a user adds personal feelings or changes task objectives, prompting instructions to be replanned. Lastly, an additional feedback and memory mechanism is adapted to adjust LNM & RLNM, incorporating behaviors from past trajectories. The real-time evaluation and feedback mechanism provides an adaptive heuristic to take advantage of. For instance, the DRL-based action execution model maintains a lower bound when the LM-based model encounters explicit mistakes or dangers, thus providing fundamental performance as the social navigation benchmark. On the other hand, the LM-based action execution model incorporates user personal modifications from user language to improve the robustness of DRL-based action execution model.
@@
Herein, the interactive social robot navigation problem is formulated as a Dec-POSMDP (decentralized particularly observable semi-Markov decision process) problem based on [31], characterized by the tuple $\langle\mathcal{S},\mathcal{U},\mathcal{A},\Omega,\mathcal{O},\mathcal{P},\mathcal{R},\mathbf{R},\mathcal{C},\mathcal{S}_{0},\gamma,\rm N,\Upsilon\rangle$. Here, $\rm\mathbf{{s}}_{\rm{t}}=[\rm\mathbf{s}^{r}_{\rm{t}},\rm\mathbf{s}^{h_{1}}_{\rm{t}},\cdots,\rm\mathbf{s}^{h_{N}}_{\rm{t}}]\in\mathcal{S}$ denotes the fully observable and unobserved states of the robot and humans at the $\rm t$-th timestep, belong to the state space, with the observable state denoted by ${\rm\mathbf{s}^{o}_{\rm{t}}}=[p_{\rm x},p_{\rm y},v_{\rm{x}},v_{\rm{y}},\rho]$ covering individual position, velocity, and radius information that can be estimated by the robot. Accordingly, pedestrians’ personal preferences and intent goals remain unobserved by robots, represented as $\mathbf{s}^{\rm{uo}}_{\rm{t}}=[{g}_{\rm{x}},{g}_{\rm{y}},v_{\rm{pref}}]$. Moreover, $\mathrm{{u}_{t}}\in\mathcal{U}$ represents robot macro-action (MA), such as waypoint locations or robot operations, which are adaptive to real-time user requirements or feedback, while robot local-action (LA) are denoted by $\mathrm{a_{t}}=[v_{x},v_{y}]\subseteq\mathcal{A}$, representing velocity. $\mathcal{O}\rm{(s^{o}|(s,a))}$ denotes the observation probability of the robot in the observation space $\Omega$, and $\mathcal{P}$ represents the state transition probability. $\mathcal{R},~{}\text{and}~{}\mathbf{R}$ represent the MA reward and LA hand-crafted reward functions separately. The MA reward is generated by the following objective: ${\mathcal{R}}(\mathrm{{s},{u}})=\mathop{\arg\max}\mathbb{E}[\textstyle\sum_{\rm t=0}^{\rm T}\gamma^{\rm t}{\mathbf{R}}({\rm s}_{\rm t},{\rm a}_{\rm t})|{\rm a}_{\rm t}\sim{\rm u}]$. Additionally, $\mathcal{C}$ represents the conditional function, which can be updated by user language comments from LLM, $\mathcal{S}_{0}$ is the initial distribution, $\rm N$ is the number of pedestrians, $\gamma\in[0,1]$ is a discount factor, and $\Upsilon$ denotes the LLM global guidance wherein task objectives, user preferences, and robot properties are involved.
@@
The interactive social navigation task statement can be viewed as a condition of a single robot from the multi-social robot navigation task definition [31]. The simulation environment also adheres to the same robot kinematic and dynamic configurations [8, 6]. For further definitions and theorems of Dec-POSMDP, refer to [33, 31].